OpenStudy (anonymous):

Is there a closed form for the series \[\sum_{n\ge1}\frac{(-1)^{n!}}{n}\,?\]

OpenStudy (anonymous):

Heck, does it even converge?

OpenStudy (empty):

I think it diverges because after n=2 we lose the interesting aspect of \( (-1)^{n!} = 1\) So it simplifies to basically the harmonic series.

OpenStudy (anonymous):

Oh, haha didn't notice that \(n!\) would be even. Never mind...

OpenStudy (dan815):

you're forgiven

OpenStudy (anonymous):

Thanks dan :P Well I suppose a more general question might be in order. Something like, what conditions on \(b_n\) are sufficient such that \[\sum_{n\ge1}\frac{(-1)^{b_n}}{n}\] converges?

OpenStudy (empty):

Yessssssssssssss

OpenStudy (dan815):

how about \[\sum_{n\ge1}\frac{(-1)^{-1^{n}}}{n}\]

OpenStudy (dan815):

lmao that is actually such a useless series now that i think about it

OpenStudy (empty):

You know what this is really quite fascinating @SithsAndGiggles I think we already know if it alternates it converges. But what about if we have: \(b_{3n} = 1\) \(b_{3n+1} = 0\) \(b_{3n+2}=0\) we would have the series: \[1 + \frac{1}{2} - \frac{1}{3} + \frac{1}{4} + \frac{1}{5} - \frac{1}{6} + \cdots \] Does this converge? :O

OpenStudy (anonymous):

So two positive, one negative, is that the general pattern?

OpenStudy (empty):

Yeah, exactly.

OpenStudy (dan815):

i want to say it converges

OpenStudy (empty):

Or I could try to generalize the problem like you two have done and try to make the most complicated looking but really actually trivial version of the harmonic series.

OpenStudy (empty):

:P

OpenStudy (empty):

Ok dan if it converges then can we say it will converge for any value? Let's say we only make one value negative every 1,000 numbers?

OpenStudy (dan815):

i think it will still

OpenStudy (dan815):

until infinity

OpenStudy (dan815):

it feels like its right on the edge

OpenStudy (empty):

The difference between 1 and .999... ? :P

OpenStudy (empty):

That reminds me, the p-series test as they called it was that if p > 1 this series converges: \[ \sum_{n=1}^\infty \frac{1}{n^p}\] So you may be on to something here.

OpenStudy (dan815):

ya thats what i was thinking about

OpenStudy (dan815):

hmm how about rewriting it like this

OpenStudy (dan815):

maybe it might diverge actually

OpenStudy (dan815):